An object moving with uniformly accelerated motion moves in a straight line. Its initial velocity is 10 m/s and it travels at total distance of 700 m in 30 s. Compute the object’s (a) acceleration and (b) final velocity.


1
Expert's answer
2021-10-20T10:42:49-0400

(a)

d=v0t+12at2,d=v_0t+\dfrac{1}{2}at^2,a=2(dv0t)t2=2×(700 m10 ms×30 s)(30 s)2=0.89 ms2.a=\dfrac{2(d-v_0t)}{t^2}=\dfrac{2\times(700\ m-10\ \dfrac{m}{s}\times30\ s)}{(30\ s)^2}=0.89\ \dfrac{m}{s^2}.

(b)

v=v0+at=10 ms+0.89 ms2×30 s=36.7 ms.v=v_0+at=10\ \dfrac{m}{s}+0.89\ \dfrac{m}{s^2}\times30\ s=36.7\ \dfrac{m}{s}.